OpenStudy (anonymous):

What number must you add to complete the square? x^2 + 26x = 68

OpenStudy (gorv):

x2+2*13*x-68=0 x2+2*13*x+169-169-68=0

OpenStudy (gorv):

we need to add 169+68=237